OpenStudy (anonymous):

What is the formula for finding the area of a sector of a circle with the radius r?

OpenStudy (anonymous):

But, you have to do something that proofs that formula or just to write the formula?

OpenStudy (cwtan):

A=(1/2)r^2\[\theta\] whr \[\theta\] is in radian

OpenStudy (anonymous):

You need to also know the angle of the sector.

OpenStudy (asnaseer):

this shows you: http://www.mathwords.com/a/area_sector_circle.htm

OpenStudy (anonymous):

Thanks:)

OpenStudy (accessdenied):

The area of a full circle is \(\pi r^2\). If we divide it into sectors, we would have a fraction of that area with respect to how large the angle is at the vertex of that sector. If we use degrees, there are 360 degrees in a circle. If we use radians, there are \(2\pi\) radians. Degrees: \(\large \frac{d}{360} \pi r^2 \quad d = \text{degrees}\) Radians: \(\large \frac{\theta}{2\pi} \pi r^2 = \frac{\theta}{2} r^2 \quad \theta = \text{radians}\)